# Marginal treatment effects in the absence of instrumental variables

### v1

@@ 4 Estimation
Our identification strategy for the MTE implies a separate estimation procedure that works with the partially linear regression functions (5) by treatment status. Specifically, the pairwise difference estimator for semiparametric selection models (Ahn and Powell, 1993) is recommended due to its computational simplicity and well-established asymptotic properties. Suppose that $\left\{\left(Y_{i},D_{i},X_{i}\right):i=1,2,\cdots,n\right\}$ is a random sample of observations on $\left(Y,D,X\right)$. In the first step, we estimate the nonparametrically specified propensity score using the kernel method, that is,
@@
where $h_{1l}$, $l=1,2,\cdots,\dim\left(X^{C}\right)$, are bandwidths and $k_{1}$ is a univariate kernel function. If the dimension of $X^{D}$ is large, then a smoothed kernel for discrete covariates (Racine and Li, 2004) can be applied as a substitute for the indicator function, to alleviate the potential problem of inadequate observations in each data cell divided by the support of $X^{D}$. If the number of continuous covariates is not small either, then the well-known curse of dimensionality will appear, and a linear-index specification may thus be practically more relevant when modeling the propensity score. The index should include a series of interaction terms and quadratic or even higher-order terms of the continuous covariates to supply sufficient nonlinear variation necessary for the identification. The linear-index propensity score can be estimated by parametric probit/logit or semiparametric methods (e.g., Powell et al., 1989; Ichimura, 1993; Klein and Spady, 1993; Lewbel, 2000), depending on the distributional assumption on the error term.
In the second step, we estimate the linear coefficients $\beta_{d}$ for each $d$ through a weighted pairwise-difference least squares regression:
@@
The pairwise difference estimator has the advantage of having a closed-form expression, so we need not solve any formidable optimization problems. However, it faces the challenging problem of bandwidth selection, similar to most semiparametric estimation methods. Alternatively, we can consider imposing a parametric specification on the unobservable heterogeneity of the MTE such that $E\left[\left.U_{d}\right|V=v\right]=E\left[\left.U_{d}\right|V=v;\theta_{d}\right]$ for finite dimensional $\theta_{d}$, e.g., the polynomial specification that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}v^{j}$ or the normal polynomial specification that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}\Phi^{-j}\left(v\right)$. In the latter, setting $J=1$ will match Heckman’s normal sample selection model. Under the parametric restriction, the second step becomes a global regression $E\left[Y\left|X,D=d\right.\right]=X^{\prime}\beta_{d}+g_{d}\left(P;\theta_{d}\right)$ with parameterized selection bias correction term $g_{d}\left(p;\theta_{d}\right)$ for each $d$; therefore, the tuning of $h_{2}$ and $h_{3}$ is circumvented. Another advantage of a parametrically specified second step is the lower computational burden compared with that of the pairwise difference estimator defined by double summation requiring a squared amount of calculation (Pan and Xie, 2023).
Notably, the local IV (LIV) estimation procedure can be adapted to our model, though it needs no IV. Unlike the separate estimation procedure, the adapted LIV approach is based on whole-sample regression:
@@
$=$ $E\left[Y_{0}\left|X\right.\right]+E\left[\left(Y_{1}-Y_{0}\right)\left|X,D=1\right.\right]\Pr\left(D=1\left|X\right.\right)$
@@
estimating $\beta_{d}$ for $d=0,1$ and functions $q$ and $q^{\left(1\right)}$ would be sufficient. Given the estimated propensity score, we can likewise use the pairwise difference principle to obtain
@@
$\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\hat{\delta}+\hat{q}^{\left(1\right)}\left(v\right),$ (16)
@@
This adapted LIV approach is computationally more convenient than the separate estimation procedure. Furthermore, if we accept a parametric specification on $E\left[U_{1}-U_{0}\left|V=v\right.\right]$ and thus on $q\left(p\right)$, the adapted LIV approach can be implemented through parametric least squares regression.
In summary, identification based on functional form can accommodate most of the frequently-used estimation procedures in a typical IV model.
@@ 5 Empirical application
We use the data provided by De Haan and Leuven (2020), which are from Round 16 (1994 survey year) of the National Longitudinal Study of Youth 1979 (NLSY79). The sample is restricted to the 1960–64 cohorts, because the first cohort eligible for Head Start was born in 1960. In addition, the sample excludes individuals who participated in any preschool programs other than Head Start, implying that we estimate the returns to Head Start relative to informal care. The treatment variable is whether the respondents attended the Head Start program as a child, and the outcome variables are the respondents’ highest years of education and logarithmic yearly wage incomes in their early 30s (they were between 30 and 34 years old in 1994). The covariates are age, gender, race, parental education, and family income in 1978. We refer the reader to the original paper for additional details on the data, sample, and variables. For our analysis, since we apply nonparametric estimation in the first step, we recode parental education into two categories to reduce the number of the data cells or subsamples split by different values of the discrete covariates. Specifically, we redefine parental education as a binary variable equal to one if at least one parent went to college or zero if both parents are high school graduates or lower.
In the first step, we nonparametrically regress the treatment variable (Head Start) on the covariates to generate propensity scores for the sample, by employing the kernel estimation method in (8) and (9) with the rule-of-thumb bandwidth. Figure 1 plots the frequency distribution of the estimated propensity score by treatment status. The figure shows that the propensity score in our sample follows a bimodal distribution, with the main peak being at approximately 0.5 for the participants and approximately 0.1 for the nonparticipants. To reduce the potential impact of outliers, we trim the observations of the 1% smallest and 1% largest propensity scores in the second-step estimation. This trimming leads to a common support ranging from 0 to 0.6, as indicated by the two dashed vertical lines.
@@
Note that only one continuous covariate exists in our data, that is, family income in 1978. Thus, we invoke Assumption NL1, which requires the propensity score to be a nonmonotonic function of the continuous covariate, given certain values of the discrete covariates. To verify Assumption NL1, we estimate the propensity score as a univariate function of family income in 1978 by using the largest data cell (32 years old, male, white race, parental education being high school or lower), with 175 observations, and plot it in Figure 2. We find that though low family income is an important eligibility criterion for enrollment in Head Start, the probability of participation is not simply a monotonically decreasing function of family income, possibly owing to the parents’ self-selection. Hence, Assumption NL1 is fulfilled.
@@
Given the estimated propensity score, we then estimate the MTE and summary treatment effect measures, such as the ATE, TT, and TUT, in the second step. We adopt the separate estimation procedure, since it generally exploits more identifying information behind the data than the LIV procedure (Brinch et al., 2017). Figure 3 shows the estimation results under Heckman’s normal specification that $E\left[\left.U_{d}\right|V=v\right]=\rho_{d}\Phi^{-1}\left(v\right)$ for $d=0,1$, where $\Phi\left(\cdot\right)$ is the CDF of the normal distribution, which can be derived by assuming that $\left(U_{d},V\right)$ follows a bivariate normal distribution with covariance $\rho_{d}$. The MTE curves, evaluated at the mean values of $X$, relate the unobserved component $U_{1}-U_{0}$ of the treatment effect to the unobserved component $V$ of the treatment choice. A high value of $V$ implies a low probability of treatment; thus, we interpret $V$ as resistance to participation in Head Start. The MTE curve for wage income plotted in panel B decreases with resistance, revealing a pattern of selection on gains, as expected. In other words, based on the unobserved characteristics, the children who were most likely to enroll in Head Start benefitted the most from the program in terms of their labor market outcome. However, when the outcome is educational attainment in panel A, the positive slope of the MTE curve points to a pattern of reverse selection on gains. As a consequence, the TUT exceeds the ATE, which in turn exceeds the TT. The same pattern was observed by Cornelissen et al. (2018) when estimating school-readiness return to a preschool program in Germany. This phenomenon may be attributed partly to the fact that parents have their own objectives in deciding childcare arrangements. Nevertheless, the disagreement between selection patterns for the two outcomes raises concern about the possible functional form misspecification of the normal MTE, which is strictly restricted to be monotonic in the resistance to treatment. Therefore, we consider a nonparametric specification for $E\left[\left.U_{d}\right|V=v\right]$ and implement a semiparametric separate estimation procedure.
@@
Based on $\hat{\beta}_{1}$ and $\hat{\beta}_{0}$ reported in Table 1, and the separate estimates of $E\left[\left.U_{d}\right|V=v\right]$ for $d=0,1$ under the semiparametric specification, we estimate the marginal structural functions
@@
where the last equality follows from the independence of $X$ and $V$ and from the conditional mean independence of $X$ and $U_{d}$ given $V$. Plugging in proper estimates of each component in the above equations, we obtain the estimated marginal response functions and plot them in Figure 6. In particular, the marginal observed outcome curves relate the individuals with significant positive returns (with medium $V$ in Panel A and small $V$ in Panel B) to those with low education and low wage income, thereby bridging our findings on the MTE and those of De Haan and Leuven (2020) on the distributional treatment effect.
@@ 6 Conclusion
We propose a novel method for defining, identifying, and estimating the MTE in the absence of IVs. Our MTE model allows all the covariates to be correlated to the structural treatment error and does not require IVs. In this model, we define the MTE based on a reduced-form treatment error that (i) is uniformly distributed on the unit interval, (ii) is statistically independent of all the covariates, and (iii) has several economic meanings. The independence property facilitates the identification of our defined MTE. We provide sufficient conditions under which the MTE can be point identified based on functional form. The conditions are standard in a certain sense. The linearity and nonlinearity assumptions are the foundation of identification based on functional form, and can make sense to most empirical studies. The conditional mean independence assumption is equivalent to the separability assumption commonly imposed in the MTE literature, and is implied by and much weaker than the full independence assumption. We prove the identification by using a construction method. Our identification strategy allows the adaptation of most of the existing estimation procedures for conventional IV-MTE models, such as separate estimation, local IV estimation, parametric estimation, and semiparametric estimation. For the empirical application, we evaluate the MTE of the Head Start program on long-term education and labor market outcomes, in which an IV for Head Start participation is difficult to acquire. We find significant positive effects for the individuals with medium-level or low resistance to treatment, but substantial heterogeneity exists.
Appendix
A.1 Proof of Theorem 1
@@
When $\dim\left(X^{C}\right)=1$, Assumption NL1 implies that
Hence, $\beta_{d}^{C}$ is identified as
When $\dim\left(X^{C}\right)\geq 2$, for $k,j\in\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ satisfying Assumption NL2, taking the partial derivatives of $m_{d0}\left(x^{C}\right)$ with respect to $x_{k}^{C}$ and $x_{j}^{C}$ yields that
@@
It follows from Assumption NL2 (i)-(ii) that
@@
which is linear in $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. The same equation is obtained if we evaluate the expression at another point $\tilde{x}^{C}$ that satisfies Assumption NL2 (iii)-(iv), which gives
@@
Assumption NL2 (v) ensures that the determinant of $\Psi$ is nonzero, which shows that $\Psi$ is nonsingular. Therefore, equation (A.3) can be solved for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$ by inverting $\Psi$, thereby identifying $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Given identification of $\beta_{d,k}^{C}$, we can then identify all other coefficient $\beta_{d,l}^{C}$ in $\beta_{d}^{C}$ by solving (A.2) with the subscript $j$ replaced by $l$, which gives
@@
A.2 Further discussion on the nonlinearity assumption
In light of the proof of Theorem 1, the local irrelevance assumption NL1 can be replaced by a differential version when there is only one continuous covariate, as given in the following theorem.
Theorem A.1.
Theorem 1 holds if Assumption NL1 is replaced by that a constant $x^{C}$ exists in the support of $X^{C}$ such that functions $\pi_{0}$ and $m_{d0}$ are differentiable at $x^{C}$ and $\left.\pi_{0}^{\left(1\right)}\left(x^{C}\right)=d\pi_{0}\left(x^{C}\right)\right/dx^{C}=0$.

### v2

@@ 4 Estimation
Our identification strategy for the MTE implies a separate estimation procedure that works with the partially linear regression (7) by each treatment status. In particular, we recommend the kernel-weighted pairwise difference estimation method proposed by Ahn and Powell (1993) because of its computational simplicity, well-established asymptotic properties, and, most importantly, its capacity to effectively leverage the overidentifying information that underlies the data. Suppose that $\left\{\left(Y_{i},D_{i},X_{i}\right):i=1,2,\cdots,n\right\}$ is a random sample of observations on $\left(Y,D,X\right)$. In the first step, we estimate the nonparametrically specified propensity score using the kernel method, that is,
@@
where $h_{1l}$, $l=1,2,\cdots,\dim\left(X^{C}\right)$, are bandwidths and $k_{1}$ is a univariate kernel function. If the dimension of $X^{D}$ is large, then a smoothed kernel for discrete covariates (Racine and Li, 2004) can be applied as a substitute for the indicator function, to alleviate the potential problem of inadequate observations in each data cell divided by the support of $X^{D}$. If the number of continuous covariates is not small either, then the well-known curse of dimensionality will appear, and a linear-index specification may thus be practically more relevant when modeling the propensity score. The index should include a series of interaction terms and quadratic or even higher-order terms of continuous covariates to supply sufficient nonlinear variation. The linear-index propensity score can be estimated by parametric probit/logit or semiparametric methods (e.g., Powell et al., 1989; Ichimura, 1993; Klein and Spady, 1993; Lewbel, 2000), depending on the distributional assumption on the error term.
In the second step, we estimate $\beta_{d}$ for each $d$ through a weighted pairwise difference least squares regression:
@@
The kernel-weighted pairwise difference estimator has the advantage of having a closed-form expression, so we need not solve any formidable optimization problems. However, it faces the challenging problem of bandwidth selection, similar to most semiparametric estimation methods. Alternatively, we can consider imposing a parametric specification on the unobservable heterogeneity of the MTE such that $E\left[\left.U_{d}\right|V=v\right]=E\left[\left.U_{d}\right|V=v;\theta_{d}\right]$ for finite dimensional $\theta_{d}$, e.g., the polynomial specification that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}v^{j}$ or the normal polynomial specification that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}\Phi^{-j}\left(v\right)$. In the latter, setting $J=1$ will match Heckman’s normal sample selection model. Under the parametric restriction, the second step becomes a global regression $E\left[Y\left|X,D=d\right.\right]=X^{\prime}\beta_{d}+g_{d}\left(P;\theta_{d}\right)$ with parameterized selection bias correction term $g_{d}\left(p;\theta_{d}\right)$ for each $d$; therefore, the tuning of $h_{2}$ and $h_{3}$ is circumvented. Another advantage of a parametrically specified second step is the lower computational burden compared with that of the pairwise difference estimator defined by double summation which requires a squared amount of calculation (Pan and Xie, 2023).
Notably, the local IV (LIV) estimation procedure can be adapted to our model as well, though it needs no IV. Unlike the separate estimation procedure, the adapted LIV approach is based on a whole-sample regression:
@@
$=$ $E\left[Y_{0}\left|X\right.\right]+E\left[\left.Y_{1}-Y_{0}\right|X,D=1\right]\Pr\left(D=1\left|X\right.\right)$
@@
the MTE is equal to the derivative of the regression function (13) with respect to $P$. As a consequence, it would be sufficient to estimate $\beta_{d}$ for $d=0,1$ and functions $q$ and $q^{\left(1\right)}$. Given the estimated propensity score, we can likewise use the pairwise difference principle to obtain
@@
$\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\hat{\delta}+\hat{q}^{\left(1\right)}\left(v\right),$
@@
This adapted LIV approach is computationally more convenient than the separate estimation procedure. Furthermore, if we accept a parametric specification on $E\left[U_{1}-U_{0}\left|V=v\right.\right]$ and thus on $q\left(p\right)$, the adapted LIV approach can be implemented through a parametric least squares regression.
In summary, the identification based on functional form can accommodate most of the frequently-used estimation procedures in a typical IV model.
@@ 5 Empirical application
We use the data provided by De Haan and Leuven (2020), which are from Round 16 (1994 survey year) of the National Longitudinal Study of Youth 1979 (NLSY79). The sample is restricted to the 1960–64 cohorts, because the first cohort eligible for Head Start was born in 1960. In addition, the sample excludes individuals who participated in any preschool programs other than Head Start, implying that we estimate the returns to Head Start relative to informal care. The treatment variable is whether the respondents attended the Head Start program as a child, and the outcome variables are the respondents’ highest years of education and logarithmic yearly wage incomes in their early 30s (they were between 30 and 34 years old in 1994). The covariates are age, gender, race, parental education, and family income in 1978. We refer the reader to the original paper for additional details on the data, sample, and variables. For our analysis, since we apply nonparametric estimation in the first step, we recode parental education into two categories to reduce the number of the data cells or subsamples split by different values of discrete covariates. Specifically, we redefine parental education as a binary variable equal to one if at least one parent went to college or zero if both parents are high school graduates or lower.
@@
We first verify the credibility of Assumptions S and NL for the Head Start data. Table 1 lists the support of the nonparametrically estimated propensity score for all data cells that are split by different values of discrete covariates. We find that the 34th data cell has a full support on the unit interval. Therefore, Assumption S must hold if we choose the corresponding values (32 years old, female, black race, parental education being college or higher) as the benchmark. We then verify Assumption NL for the 34th data cell. Note that only one continuous covariate exists in the data, that is, family income in 1978. Hence, we invoke Assumption NL1, which requires the propensity score function to have at least one stationary point. The left panel of Figure 1 plots the estimated propensity score as a univariate function of the continuous covariate for the 34th data cell. We find that the propensity score is highly nonlinear and nonmonotonic in family income in 1978, possibly owing to the parents’ various self-selection on participation in the program. So Assumption NL1 is fulfilled.
Furthermore, the largest data cell (i.e., the 29th data cell with 175 observations) can be considered an alternative candidate for the benchmark (32 years old, male, white race, parental education being high school or lower). Nevertheless, closer examination of Assumption S is necessary for this data cell, as the support of the propensity score is now limited to $\left[0,0.194\right]$, which is totally separate from the supports for some other data cells. To this end, we need to find a distinct value of each discrete variable such that the propensity score has overlapping support with $\left[0,0.194\right]$. For example, if the value of age is altered from 32 to 30 (or 31, 33, 34) while the other discrete variables remain unaltered, we will go to the 5th (or 17th, 41st, 53rd) data cell with support $\left[0,0.177\right]$ (or $\left[0,0.242\right],\left[0,0.154\right],\left[0,0.204\right]$), which is overlapped with $\left[0,0.194\right]$ as required. Altering the values of gender, race, and parental education leads to the 35th, 25th (or 27th), and 30th data cells with supports $\left[0,0.526\right]$, $\left[0,0.771\right]$ (or $\left[0,0.623\right]$), and $\left[0,0.061\right]$, respectively, which all meet the overlapping condition as well. Therefore, Assumption S holds for the largest data cell. The validation of Assumption NL1 for this data cell can be demonstrated as well by the existence of stationary points of the estimated propensity score function, as illustrated in the right panel of Figure 1.
We next turn to the estimation. In the first step, we nonparametrically regress the treatment variable (Head Start) on all of the covariates to generate propensity scores for the sample, by employing the kernel estimation method in (10) and (11) with the rule-of-thumb bandwidth. Figure 2 plots the frequency distribution of the estimated propensity score by treatment status. The figure shows that the propensity score in our sample follows a bimodal distribution, with the main peak being at approximately 0.5 for the participants and approximately 0.1 for the nonparticipants. To reduce the potential impact of outliers, we trim the observations of the 1% smallest and 1% largest propensity scores in the following steps. This trimming leads to a common support ranging from 0 to 0.6, as indicated by the two dashed vertical lines. Given the estimated propensity score, we then estimate in sequence the linear coeffecients, MTE, and summary treatment effect measures including ATE, TT, and TUT. We adopt the separate estimation procedure, since it generally exploits more identifying information behind the data than the LIV procedure (Brinch et al., 2017).
@@
Figure 3 shows the estimated MTE under Heckman’s normal specification that $E\left[\left.U_{d}\right|V=v\right]=\rho_{d}\Phi^{-1}\left(v\right)$ for $d=0,1$, where $\Phi\left(\cdot\right)$ is the CDF of normal distribution, which can be derived by assuming that $\left(U_{d},V\right)$ follows a bivariate normal distribution with covariance $\rho_{d}$. The MTE curves, evaluated at the mean values of $X$, relate the unobserved component $U_{1}-U_{0}$ of the treatment effect to the unobserved component $V$ of the treatment choice. A high value of $V$ implies a low probability of treatment; thus, we interpret $V$ as resistance to participation in Head Start. The MTE curve for wage income plotted in panel B decreases with resistance, revealing a pattern of selection on gains, as expected. In other words, based on the unobserved characteristics, the children who were most likely to enroll in Head Start benefitted the most from the program in terms of their labor market outcome. However, when the outcome is educational attainment in panel A, the positive slope of the MTE curve points to a pattern of reverse selection on gains. In consequence, the TUT exceeds the ATE, which in turn exceeds the TT. The same pattern was observed by Cornelissen et al. (2018) when estimating school-readiness return to a preschool program in Germany. This phenomenon may be attributed partly to the fact that parents have their own objectives in deciding childcare arrangements. Nevertheless, the disagreement between selection patterns for the two outcomes raises concern about the possible functional form misspecification of the normal MTE, which is strictly restricted to be monotonic in the resistance to treatment. Therefore, we consider a nonparametric specification for $E\left[\left.U_{d}\right|V=v\right]$ and implement a semiparametric separate estimation procedure.
@@
Based on $\hat{\beta}_{1}$ and $\hat{\beta}_{0}$ reported in Table 2, and the separate estimates of $E\left[\left.U_{d}\right|V=v\right]$ for $d=0,1$ under the semiparametric specification, we can estimate the marginal structural functions
@@
where the last equality follows from the independence of $X$ and $V$ and from the conditional mean independence of $X$ and $U_{d}$ given $V$. Plugging in proper estimates of each component in the above equations, we obtain the estimated marginal response functions and plot them in Figure 6. In particular, the marginal observed outcome curves relate the individuals with significant positive returns (with medium $V$ in Panel A and small $V$ in Panel B) to those with low education and low wage income, thereby bridging our results on the MTE and those of De Haan and Leuven (2020) on the distributional treatment effect.
@@ 6 Conclusion
We propose a novel method for defining, identifying, and estimating the MTE in the absence of IVs. Our MTE model allows all the covariates to be correlated to the structural treatment error. In this model, we define the MTE based on a reduced-form treatment error that (i) is uniformly distributed on the unit interval, (ii) is statistically independent of covariates, and (iii) has several economic meanings. The independence property facilitates the identification of our defined MTE. We provide sufficient conditions under which the MTE can be point identified based on functional form. The conditions are standard in a certain sense. The conditional mean independence assumption is equivalent to the separability assumption commonly imposed in the MTE literature, and is implied by and much weaker than the full independence assumption. The linearity and nonlinearity assumptions are the foundation of identification based on functional form, and can make sense to most empirical studies. We prove the identification by using a construction method. Our identification strategy allows the adaptation of most of the existing estimation procedures for conventional IV-MTE models, such as separate estimation, LIV estimation, parametric estimation, and semiparametric estimation. For the empirical application, we evaluate the MTE of the Head Start program on long-term education and labor market outcomes, in which an IV for Head Start participation is difficult to acquire. We find significant positive effects for individuals with medium-level or low resistance to treatment, and substantial heterogeneity exists.
Appendix A Proof of Theorem 1
@@
When $\dim\left(X^{C}\right)=1$, by differentiating both sides of (A.1) at $\tilde{x}^{C}$ and invoking Assumption NL1, we obtain
$m_{d0}^{\left(1\right)}\left(\tilde{x}^{C}\right)=\beta_{d}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\tilde{x}^{C}\right)\right)\pi_{0}^{\left(1\right)}\left(\tilde{x}^{C}\right)=\beta_{d}^{C},$
which identifies $\beta_{d}^{C}$ for $d=0,1$. When $\dim\left(X^{C}\right)\geq 2$, for $k,j\in\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ satisfying Assumption NL2, taking the partial derivatives of $m_{d0}\left(x^{C}\right)$ with respect to $x_{k}^{C}$ and $x_{j}^{C}$ yields that
@@
It follows from Assumption NL2.(i)-(ii) that
@@
which is linear in $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. The same equation is obtained if we evaluate the expression at another point $\tilde{x}^{C}$ that satisfies Assumption NL2.(iii)-(iv), which gives
@@
Assumption NL2.(v) ensures that the determinant of $\Psi$ is nonzero, which shows that $\Psi$ is nonsingular. Therefore, equation (A.3) can be solved for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$ by inverting $\Psi$, thereby identifying $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Given identification of $\beta_{d,k}^{C}$, we can then identify all other coefficient $\beta_{d,l}^{C}$ in $\beta_{d}^{C}$ by solving (A.2) with the subscript $j$ replaced by $l$, which gives
@@
Appendix B Further discussion on the nonlinearity assumption
In light of the proof of Theorem 1, our identification strategy may work even when no continuous covariate is available in the data. In that case, we put a discretely distributed covariate in $X^{C}$.
Theorem B.1.
Theorem 1 holds if Assumption NL1 is replaced by that there exist two different constants $x^{C}$, $\tilde{x}^{C}$ in the support of $X^{C}$ such that $\pi_{0}\left(x^{C}\right)=\pi_{0}\left(\tilde{x}^{C}\right)$.